OpenStudy (anonymous):

how do u factor this question: 2w(y-2z)-x(y-2z)

OpenStudy (anonymous):

(y-2z)(2w-x)

OpenStudy (anonymous):

It is (y-2z) (2w - x) this is by taking the common factor (y-2z) from both

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

by taking the common factor

OpenStudy (anonymous):

They're both multiplied by the common factor

OpenStudy (anonymous):

can u show me step by step, im confused on factoring equations

OpenStudy (anonymous):

Here we use the distributive law in reverse

OpenStudy (anonymous):

distributive law says a (b+d) = ab + ad This can also be put as ab + ad = a (b+d) Understood now??

OpenStudy (anonymous):

yes thank u

OpenStudy (anonymous):

but how to do u get the answer since the question has ( )

OpenStudy (anonymous):

when u have more than one number i.e. it is two terms then u have to put brackets so in your case a(b) + a(d) = a (b+d)

OpenStudy (anonymous):

oh ok, thanks again

OpenStudy (anonymous):

Since (y-2z) is common to both treat it as a and 2w as b and -x as d

OpenStudy (anonymous):

and how about x^2-2x+3x-6 how do u do that one

OpenStudy (anonymous):

Here in the first two terms x^2 and -2x we have common factor x so they can be combined using distributive law as x(x -2) Now we have to resolve the last two terms so that we get (x-2) in the bracket this can be done as follows +3 (x-2)

OpenStudy (anonymous):

So the question become x(x-2) +3 (x-2) We find (x-2) is common so it becomes (x-2) (x+3) Got it??

OpenStudy (anonymous):

oh ok, thanks

OpenStudy (anonymous):

Where r u located sanasaj???

OpenStudy (anonymous):

y

OpenStudy (anonymous):

just curious, I am frm Delhi, india

OpenStudy (anonymous):

oh ok, im from the US

OpenStudy (anonymous):

U going to school or home-study??

OpenStudy (anonymous):

home-study

OpenStudy (anonymous):

I think u shud buy books by R. D. Sharma - an Indian author - for maths. I think u must b in 7th or 8th standard